OpenStudy (anonymous):

The measure of each exterior angle of a regular polygon is 40°. How many sides does the polygon have? 12 18 20 9

OpenStudy (anonymous):

cmo i know u guys are smart plz help lol

OpenStudy (anonymous):

40 +y - 40 = 180 -40 y = 140 (sum of interior angles) use polygon formula : y=180*(n-2)/n plug in y 140=180-(360/n) -40=-360/n 40=360/n n(40)=(360/n)n 40n=360 n=9

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

have i ever told you how amazinggg you are!!! i hope u like gold!!!

OpenStudy (anonymous):

hahaaha :) thanks. and why yes, i ♥ gold